OpenStudy (anonymous):

Solve 5^(x+5)=9^x a)x≈–3.63 b)x≈–0.28 c)x≈0.07 d)x≈13.69

OpenStudy (anonymous):

i know it is not A

OpenStudy (tkhunny):

Have you considered the logarithm function?

OpenStudy (anonymous):

could you please state the function, i'm a litle rusty at algebra

OpenStudy (tkhunny):

Why are you given this problem if you've no logarithm? Is this a placement exam? If it is a placement exam, you should get this problem wrong so that the examiner will know that you need more work in this area.

OpenStudy (anonymous):

no this is review packet for finals and i forgot how to do this type of problem

OpenStudy (tkhunny):

If you say so. One for free. \(5^{x+5}=9^{x}\) Introduce the logarithm \(\log\left(5^{x+5}\right)=\log\left(9^{x}\right)\) logarithm properties \((x+5)\cdot \log(5)=x\cdot \log(9)\) Distributive property \(x\cdot \log(5) + 5\cdot \log(5)=x\cdot \log(9)\) Can you solve this for 'x'?

OpenStudy (anonymous):

so \[.698970004336019(x+5)=.9542509439325(x)\]

OpenStudy (anonymous):

so the answer is d) 13.69

OpenStudy (tkhunny):

That's kind of a mess, but you did find the right answer. I would suggest you review this section quite a bit more and brush up on your logarithms and your basic algebra. you switched to your calculator WAY too quickly.